### v2

@@ 5 Conclusion
To improve the condition number of the matrix corresponding to the operator inference linear system, future work could investigate the exploitation of nestedness to solve subproblems sequentially [2] or different choices of rank-ensuring ROM states, for example based on Chebychev points [26].
@@ A.1 Outline of proof
The special case of degree sets without gaps, (i.e. $\mathcal{I}=\mathbb{N}_{l}$), is covered by an existing result on existence and uniqueness of an interpolating multivariate polynomial [17] described in Section A.2. In Section A.3, we use this result to prove Theorem A for the special case that $\mathcal{I}$ contains only one element. I.e., we can find for all $i\in\mathcal{I}$ a homogeneous polynomial $p_{i}\in\mathcal{P}^{n}_{\{i\}}$ that interpolates in the nodes in $\mathcal{X}^{i}$. We use this result as base case in a proof by induction in Section A.5 to prove the general case where $\mathcal{I}$ contains multiple elements. There, the induction hypothesis states that Theorem A holds for a fixed degree set ${\mathcal{I}^{-}}$, i.e., there exists a unique polynomial $p^{-}$ that interpolates in the interpolation nodes in $\cup_{i\in{\mathcal{I}^{-}}}\mathcal{X}^{i}$. In the induction step, we show that Theorem A also holds for the new degree set :=$\mathcal{I}:=\{{i^{*}}\}\cup{\mathcal{I}^{-}}$ for any ${i^{*}}\in\mathbb{N}_{0}$ with ${i^{*}}<\min{\mathcal{I}^{-}}$. For this purpose, we add a new polynomial $p$ that has two key properties. First, $p$ attains suitable values in the new interpolation nodes in $\mathcal{X}^{{i^{*}}}$ such that the resulting polynomial $p^{-}+p$ attains the prescribed interpolation values $b_{j}$. Second, $p$ attains 0 in the old interpolation nodes in $\cup_{i\in{\mathcal{I}^{-}}}\mathcal{X}^{i}$ to not interfere with the interpolation property of $p^{-}$ in these nodes. In Section A.4, we show that such a polynomial $p$ exists in $\mathcal{P}^{n}_{\mathcal{I}}$.
@@ A.2 Multivariate polynomial interpolation without gaps
For any ${m}\in\mathbb{N}_{0}$, let ${\bm{X}}_{0},{\bm{X}}_{1},\dots,{\bm{X}}_{m}\in\mathbb{R}^{m}$ be the vertices of a non-degenerate ${m}$-dimensional simplex in $\mathbb{R}^{m}$. For any $l\in\mathbb{N}_{0}$, we define the lattice
(40) $B(l,{m}):=\left\{{\bm{x}}\in\mathbb{R}^{m}\Bigg{|}{\bm{x}}=\sum_{i=0}^{{m}}\lambda_{i}{\bm{X}}_{i},\;\lambda_{i}\in\mathbb{N}_{0},\;\sum_{i=0}^{{m}}\lambda_{i}=l\right\}.$
@@
for the interpolation nodes ${\bm{q}}_{i}$ in $B(l,{m})$ and arbitrary interpolation values $f_{i}\in\mathbb{R}$.
@@ A.3 Homogeneous multivariate polynomial interpolation
To prove this theorem, we trace it back to Theorem A.2 on multivariate polynomial interpolation without gaps. First, we note that $\mathcal{X}^{l}$ lies in an ($n-1$)-dimensional subspace of $\mathbb{R}^{n}$ and constitutes a lattice of degree $l$ and dimension $n-1$. Hence, we can construct a multivariate polynomial $p\in\mathcal{P}^{n}_{\mathbb{N}_{l}}$ that interpolates arbitrary values in $\mathcal{X}^{l}$. Then, we use the linear dependence of the nodes in $\mathcal{X}^{l}$ to reformulate $p$ into an equivalent polynomial in $\mathcal{P}^{n}_{\{l\}}$.
@@
We employ Theorem A.2 with ${m}=n-1$ and the vertices ${\bm{X}}_{0}={\bm{0}}\in\mathbb{R}^{n-1}$ and ${\bm{X}}_{j}$ the unit vectors in $\mathbb{R}^{n-1}$ for $j=1,\dots,n-1$. The theorem states that there exists a polynomial
@@
with $\hat{a}_{\hat{\bm{\alpha}}}\in\mathbb{R}$ for all multi-exponents ${\hat{\bm{\alpha}}}\in\{0,1,\dots,l\}^{n-1}$ with $|{\hat{\bm{\alpha}}}|\leq l$, such that
@@
for the prescribed interpolation values $f_{i}$ and the interpolation nodes $\hat{\bm{q}}_{i}$ in $B(l,n-1)$.
We define the linear map ${\bm{m}}$ that maps the vertices ${\bm{X}}_{j}$ to the unit vectors in $\mathbb{R}^{n}$. Concretely, ${\bm{m}}({\bm{X}}_{j})={\bm{e}}_{j}$, $j=1,\dots,n-1$ and ${\bm{m}}({\bm{X}}_{0})={\bm{e}}_{n}$. In view of the definition of $\mathcal{X}^{i}$ (13), we can rewrite the definition of $B(l,m)$ (40),
(45) $B(l,m)=\left\{\sum_{j=1}^{l}{\bm{a}}_{j}\,\Bigg{|}{\bm{a}}_{1},\dots,{\bm{a}}_{l}\in\left\{{\bm{X}}_{0},\dots,{\bm{X}}_{m}\right\}\right\}.$
@@
(46) ${\bm{m}}(B(l,n-1))=\left\{\sum_{j=1}^{l}{\bm{a}}_{j}\,\Bigg{|}{\bm{a}}_{1},\dots,{\bm{a}}_{l}\in\left\{{\bm{m}}({\bm{X}}_{0}),\dots,{\bm{m}}({\bm{X}}_{n-1})\right\}\right\}=\mathcal{X}^{l}.$
@@
(50) $p({\bm{x}})=\hat{p}({\bm{m}}^{-1}({\bm{x}}))=\sum_{|{\hat{\bm{\alpha}}}|\leq l}\hat{a}_{\hat{\bm{\alpha}}}({\bm{m}}^{-1}({\bm{x}}))^{\hat{\bm{\alpha}}}=\sum_{|{\hat{\bm{\alpha}}}|\leq l}\hat{a}_{\hat{\bm{\alpha}}}\prod_{i=1}^{n-1}x_{i}^{\hat{\alpha}_{i}}.$
By definition, all ${\bm{x}}\in\mathcal{X}^{l}$ satisfy
@@
with suitable $a_{\bm{\alpha}}\in\mathbb{R}$, for all ${\bm{\alpha}}\in\{0,1,\dots,l\}^{n}$ with $|{\bm{\alpha}}|=l$. Thereby, we have proven existence of a polynomial $p\in\mathcal{P}^{n}_{\{l\}}$ that satisfies (42) for any interpolation values $f_{i}$, $i=1,2,\dots,n_{l}$. Since the corresponding interpolation matrix is square, existence of a solution implies also uniqueness.
@@ A.4 Specific gappy multivariate polynomial interpolation
As explained in Section A.1, the proof of Theorem A in Section A.5 involves a polynomial $p\in\mathcal{P}^{n}_{\mathcal{I}}$ with :=$\mathcal{I}:=\{{i^{*}}\}\cup{\mathcal{I}^{-}}$ such that
@@
For any $l\in\mathbb{N}_{0}$, for any :=$\mathcal{I}:=\{{i^{*}}\}\,\cup\,{\mathcal{I}^{-}}\subset\mathbb{N}_{l}$ with ${i^{*}}<\min{\mathcal{I}^{-}}$, there exists a polynomial $p\in\mathcal{P}^{1}_{{\mathcal{I}^{-{i^{*}}}}}$ with ${\mathcal{I}^{-{i^{*}}}}:=\{i-{i^{*}}\big{|}i\in\mathcal{I}\}$ such that
@@
  • Base case: For $l=0$, $p(x)=1$ satisfies the interpolation properties for the only existing set =$\mathcal{I}=\{0\}$.
@@
  • Induction step: We consider a set $\mathcal{I}\subset\mathbb{N}_{l+1}:=\{0,1,\dots,l,l+1\}$ and the pairwise distinct, positive nodes $x_{1},\dots,x_{n_{\mathcal{I}}}\in\mathbb{R}^{+}$. If $l+1\notin\mathcal{I}$, then the induction hypothesis directly applies. Otherwise, if $l+1\in\mathcal{I}$, then the induction hypothesis holds for :=$\mathcal{I}^{-}:=\mathcal{I}\setminus\{l+1\}$ and the nodes $x_{1},\dots,x_{{n_{\mathcal{I}}}-1}$. Hence, Lemma A.2 implies that there exists a polynomial $\bar{p}\in\mathcal{P}^{1}_{\mathcal{I}^{-}}$ such that
@@
Let :=$\mathcal{I}^{-}:=\mathcal{I}\setminus\{0\}$. According to Lemma A.2 and Lemma A.4, there exists a polynomial $\bar{p}\in\mathcal{P}^{1}_{\mathcal{I}^{-}}$ such that
@@
For any $l\in\mathbb{N}$, for any ^=$\hat{\mathcal{I}}=\{{i^{*}}\}\cup{\mathcal{I}^{-}}\subset\mathbb{N}_{l}$ with ${i^{*}}<\min{\mathcal{I}^{-}}$, let =$\mathcal{I}=\{i-{i^{*}}\big{|}i\in\hat{\mathcal{I}}\}$ and $\{x_{1},\dots,x_{n_{\mathcal{I}}}\}=\hat{\mathcal{I}}$. Then Lemma A.4 implies the statement if ${i^{*}}\neq 0$ and Lemma A.6 if ${i^{*}}=0$.
@@ A.5 Gappy multivariate polynomial interpolation
  • Base case: Let =$\mathcal{I}=\{i\}$ for any $i\in\mathbb{N}_{l}$ with arbitrary $l\in\mathbb{N}_{0}$, then Theorem A.3 implies the existence and uniqueness of a polynomial $p\in\mathcal{P}^{n}_{\{i\}}$ that satisfies the interpolation property (39) in the interpolation nodes in $\mathcal{X}^{i}$.
@@
  • Induction step: We consider the degree set :=$\mathcal{I}:=\{{i^{*}}\}\cup{\mathcal{I}^{-}}$ with arbitrary ${i^{*}}\in\mathbb{N}_{l}$. Without loss of generality, we can assume ${i^{*}}<\min{\mathcal{I}^{-}}$. As described in Section A.4, we can construct a polynomial

### v3

@@ 5 Conclusion
To improve the condition number of the matrix corresponding to the operator inference linear system, future work could investigate the exploitation of nestedness to solve subproblems sequentially [2] or different choices of rank-ensuring ROM states, for example based on Chebychev points [27]. Future work should also investigate how to enable that the proposed rank-ensuring states can be specified as initial conditions in solvers that allow only a certain range of state values or require specific boundary conditions.
@@ A.1 Outline of proof
The special case of degree sets without gaps, (i.e. $\mathcal{I}=\mathbb{N}_{l}$), is covered by an existing result on existence and uniqueness of an interpolating multivariate polynomial [17] described in Section A.2. In Section A.3, we use this result to prove Theorem A for the special case that $\mathcal{I}$ contains only one element. I.e., we can find for all $i\in\mathcal{I}$ a homogeneous polynomial $p_{i}\in\mathcal{P}^{n}_{\{i\}}$ that interpolates in the nodes in $\mathcal{X}^{i}$. We use this result as base case in a proof by induction in Section A.5 to prove the general case where $\mathcal{I}$ contains multiple elements. There, the induction hypothesis states that Theorem A holds for a fixed degree set $\mathcal{I}^{-}$}, i.e., there exists a unique polynomial $p^{-}$ that interpolates in the interpolation nodes in $\cup_{i\in{\mathcal{I}^{-}}}\mathcal{X}^{i}$. In the induction step, we show that Theorem A also holds for the new degree set $\mathcal{I}:=\{{i^{*}}\}\cup{\mathcal{I}^{-}}$ for any ${i^{*}}\in\mathbb{N}_{0}$ with ${i^{*}}<\min{\mathcal{I}^{-}}$. For this purpose, we add a new polynomial $p$ that has two key properties. First, $p$ attains suitable values in the new interpolation nodes in $\mathcal{X}^{{i^{*}}}$ such that the resulting polynomial $p^{-}+p$ attains the prescribed interpolation values $b_{j}$. Second, $p$ attains 0 in the old interpolation nodes in $\cup_{i\in{\mathcal{I}^{-}}}\mathcal{X}^{i}$ to not interfere with the interpolation property of $p^{-}$ in these nodes. In Section A.4, we show that such a polynomial $p$ exists in $\mathcal{P}^{n}_{\mathcal{I}}$.
@@ A.2 Multivariate polynomial interpolation without gaps
For any ${m}\in\mathbb{N}_{0}$, let ${\bm{y}}_{0},{\bm{y}}_{1},\dots,{\bm{y}}_{m}\in\mathbb{R}^{m}$ be the vertices of a non-degenerate m{m}-dimensional simplex in $\mathbb{R}^{m}$. For any $l\in\mathbb{N}_{0}$, we define the lattice
(40) $B(l,{m}):=\left\{{\bm{x}}\in\mathbb{R}^{m}\Bigg|{\bm{x}}=\sum_{i=0}^{{m}}\lambda_{i}{\bm{y}}_{i},\;\lambda_{i}\in\mathbb{N}_{0},\;\sum_{i=0}^{{m}}\lambda_{i}=l\right\}.$
@@
for the interpolation nodes ${\bm{q}}_{i}$ in B(l,m)B(l,{m}) and arbitrary interpolation values $f_{i}\in\mathbb{R}$.
@@ A.3 Homogeneous multivariate polynomial interpolation
To prove this theorem, we trace it back to Theorem A.2 on multivariate polynomial interpolation without gaps. First, we note that $\mathcal{X}^{l}$ lies in an (n1n-1)-dimensional subspace of $\mathbb{R}^{n}$ and constitutes a lattice of degree $l$ and dimension n1n-1. Hence, we can construct a multivariate polynomial $p\in\mathcal{P}^{n}_{\mathbb{N}_{l}}$ that interpolates arbitrary values in $\mathcal{X}^{l}$. Then, we use the linear dependence of the nodes in $\mathcal{X}^{l}$ to reformulate $p$ into an equivalent polynomial in $\mathcal{P}^{n}_{\{l\}}$.
@@
We employ Theorem A.2 with m=n1{m}=n-1 and the vertices ${\bm{y}}_{0}={\bm{0}}\in\mathbb{R}^{n-1}$ and ${\bm{y}}_{j}$ the unit vectors in $\mathbb{R}^{n-1}$ for $j=1,\dots,n-1$. The theorem states that there exists a polynomial
@@
[…]
@@
for the prescribed interpolation values $f_{i}$ and the interpolation nodes $\hat{\bm{q}}_{i}$ in B(l,n1)B(l,n-1).
We define the linear map $\bm{m}$} that maps the vertices ${\bm{y}}_{j}$ to the unit vectors in $\mathbb{R}^{n}$. Concretely, ${\bm{m}}({\bm{y}}_{j})={\bm{e}}_{j}$, $j=1,\dots,n-1$ and ${\bm{m}}({\bm{y}}_{0})={\bm{e}}_{n}$. In view of the definition of $\mathcal{X}^{i}$ (13), we can rewrite the definition of $B(l,m)$ (40),
(45) $B(l,m)=\left\{\sum_{j=1}^{l}{\bm{a}}_{j}\,\Bigg|{\bm{a}}_{1},\dots,{\bm{a}}_{l}\in\left\{{\bm{y}}_{0},\dots,{\bm{y}}_{m}\right\}\right\}.$
@@
(46) ${\bm{m}}(B(l,n-1))=\left\{\sum_{j=1}^{l}{\bm{a}}_{j}\,\Bigg|{\bm{a}}_{1},\dots,{\bm{a}}_{l}\in\left\{{\bm{m}}({\bm{y}}_{0}),\dots,{\bm{m}}({\bm{y}}_{n-1})\right\}\right\}=\mathcal{X}^{l}.$
@@
(50) $p({\bm{x}})=\hat{p}({\bm{m}}^{-1}({\bm{x}}))=\sum_{|{\hat{\bm{\alpha}}}|\leq l}\hat{a}_{\hat{\bm{\alpha}}}({\bm{m}}^{-1}({\bm{x}}))^{\hat{\bm{\alpha}}}=\sum_{|{\hat{\bm{\alpha}}}|\leq l}\hat{a}_{\hat{\bm{\alpha}}}\prod_{i=1}^{n-1}x_{i}^{\hat{\alpha}_{i}}{,}$
with the entries $x_{i}$ of ${\bm{x}}\in\mathbb{R}^{n}$, $i=1,\dots,n$. By definition, all ${\bm{x}}\in\mathcal{X}^{l}$ satisfy
@@
with suitable $a_{\bm{\alpha}}\in\mathbb{R}$, for all ${\bm{\alpha}}\in\{0,1,\dots,l\}^{n}$ with $|{\bm{\alpha}}|=l$. Thereby, we have proven existence of a polynomial $p\in\mathcal{P}^{n}_{\{l\}}$ that satisfies (42) for any interpolation values $f_{i}$, $i=1,2,\dots,{{n+l-1\choose l}}$. Since the corresponding interpolation matrix is square, existence of a solution implies also uniqueness.
@@ A.4 Specific gappy multivariate polynomial interpolation
As explained in Section A.1, the proof of Theorem A in Section A.5 involves a polynomial $p\in\mathcal{P}^{n}_{\mathcal{I}}$ with $\mathcal{I}:=\{{i^{*}}\}\cup{\mathcal{I}^{-}}$ such that
@@
For any $l\in\mathbb{N}_{0}$, for any $\mathcal{I}:=\{{i^{*}}\}\,\cup\,{\mathcal{I}^{-}}\subset\mathbb{N}_{l}$ with ${i^{*}}<\min{\mathcal{I}^{-}}$, there exists a polynomial $p\in\mathcal{P}^{1}_{{\mathcal{I}^{-{i^{*}}}}}$ with ${\mathcal{I}^{-{i^{*}}}}:=\{i-{i^{*}}\big|i\in\mathcal{I}\}$ such that
@@
  • Base case: For $l=0$, $p(x)=1$ satisfies the interpolation properties for the only existing set $\mathcal{I}=\{0\}$.
@@
  • Induction step: We consider a set $\mathcal{I}\subset\mathbb{N}_{l+1}:=\{0,1,\dots,l,l+1\}$ and the pairwise distinct, positive nodes $x_{1},\dots,x_{n_{\mathcal{I}}}\in\mathbb{R}^{+}$. If $l+1\notin\mathcal{I}$, then the induction hypothesis directly applies. Otherwise, if $l+1\in\mathcal{I}$, then the induction hypothesis holds for $\mathcal{I}^{-}:=\mathcal{I}\setminus\{l+1\}$ and the nodes $x_{1},\dots,x_{{n_{\mathcal{I}}}-1}$. Hence, Lemma A.2 implies that there exists a polynomial $\bar{p}\in\mathcal{P}^{1}_{\mathcal{I}^{-}}$ such that
@@
Let $\mathcal{I}^{-}:=\mathcal{I}\setminus\{0\}$. According to Lemma A.2 and Lemma A.4, there exists a polynomial $\bar{p}\in\mathcal{P}^{1}_{\mathcal{I}^{-}}$ such that
@@
For any $l\in\mathbb{N}$, for any $\hat{\mathcal{I}}=\{{i^{*}}\}\cup{\mathcal{I}^{-}}\subset\mathbb{N}_{l}$ with ${i^{*}}<\min{\mathcal{I}^{-}}$, let $\mathcal{I}=\{i-{i^{*}}\big|i\in\hat{\mathcal{I}}\}$ and $\{x_{1},\dots,x_{n_{\mathcal{I}}}\}=\hat{\mathcal{I}}$. Then Lemma A.4 implies the statement if ${i^{*}}\neq 0$ and Lemma A.6 if ${i^{*}}=0$.
@@ A.5 Gappy multivariate polynomial interpolation
  • Base case: Let $\mathcal{I}=\{i\}$ for any $i\in\mathbb{N}_{l}$ with arbitrary $l\in\mathbb{N}_{0}$, then Theorem A.3 implies the existence and uniqueness of a polynomial $p\in\mathcal{P}^{n}_{\{i\}}$ that satisfies the interpolation property (39) in the interpolation nodes in $\mathcal{X}^{i}$.
@@
  • Induction step: We consider the degree set $\mathcal{I}:=\{{i^{*}}\}\cup{\mathcal{I}^{-}}$ with arbitrary ${i^{*}}\in\mathbb{N}_{l}$. Without loss of generality, we can assume ${i^{*}}<\min{\mathcal{I}^{-}}$. As described in Section A.4, we can construct a polynomial
@@
Appendix B Algorithm for generating ROM state set
We present two different approaches how to implement the procedure to generate the rank-ensuring ROM state sets $\mathcal{X}^{i}$ as defined in (13) in line 7 of Algorithm 1. As shown in Algorithm 2 for polynomial degree $i=3$, the procedure essentially comprises 3 nested for-loops. The first approach shown in Algorithm 3 constructs these nested for-loops for arbitrary $i$ via recursion. This recursion makes the algorithm rather non-intuitive. While this non-intuitivity might be unavoidable to handle arbitrary polynomial degrees $i$, a more intuitive approach is possible in practice. In practice, we know a priori which polynomial degrees $i$ are involved in the considered problems. For example, in the numerical experiments in Section 4, the only polynomial degrees involved are $\{1,2,3,8\}$. Consequently, we can use the hard-coded discrimination of cases shown in Algorithm 4 as alternative to Algorithm 3, with the procedures for $i=1,2,8$ implemented analogous to Algorithm 2.
Algorithm 2 Generate rank-ensuring ROM state set $\mathcal{X}^{3}$
1:procedure GenerateROMstateSet3( ROM dimension $n$ )
2:  Initialize $\mathcal{X}^{3}\leftarrow\emptyset$
3:  for $j_{1}=1,\dots,n$ do
4:   for $j_{2}=j_{1},\dots,n$ do
5:     for $j_{3}=j_{2},\dots,n$ do
6:      $\mathcal{X}^{3}\leftarrow\mathcal{X}^{3}\cup{\bm{e}}_{j_{1}}+{\bm{e}}_{j_{2}}+{\bm{e}}_{j_{3}}$
7:     end for
8:   end for
9:  end for
10:  return $\mathcal{X}^{3}$
11:end procedure
Algorithm 3 Generate rank-ensuring ROM state set $\mathcal{X}^{i}$ - recursive algorithm
1:procedure GenerateROMstateSet(
ROM dimension $n$,
polynomial degree $i\in\mathbb{N}_{0}$
2:  return recursiveROMstateSet( ROM dimension $n$, polynomial degree $i$,
——————————————————- current for loop depth ${i^{*}}=0$,
——————————————————- current index sequence $\bm{j}$}=[],
——————————————————- current ROM state set $\mathcal{X}^{i}=\emptyset$ )
3:end procedure
4:procedure recursiveROMstateSet( ROM dimension $n$, polynomial degree $i\in\mathbb{N}_{0}$,
———————————- current for loop depth $i^{*}$},
———————————- current index sequence ${\bm{j}}=[j_{1},\dots,j_{i^{*}}]$, # always has length $i^{*}$}
———————————- current ROM state set $\mathcal{X}^{i}$ )
5:  if ${i^{*}}=i$ then
6:   Initialize $\bar{\bm{x}}\leftarrow{\bm{0}}\hskip 10.00002pt\in\mathbb{R}^{n}$
7:   for k=1,…,i do
8:     $\bar{\bm{x}}\leftarrow\bar{\bm{x}}+{\bm{e}}_{j_{k}}$
9:   end for
10:   return $\mathcal{X}^{i}\cup\bar{\bm{x}}$
11:  else
12:   # Perform next inner for-loop
13:   for $j_{{i^{*}}+1}=j_{i^{*}},\dots,n$ do # with $j_{0}=1$
14:     return recursiveROMstateSet( ROM dimension $n$, polynomial degree $i$,
——————————————————- current for loop depth ${i^{*}}+1$,
——————————————————- current index sequence $j_{1},\dots,j_{{i^{*}}+1}]$,
——————————————————- current ROM state set $\mathcal{X}^{i}$ )
15:   end for
16:  end if
17:end procedure
Algorithm 4 Generate rank-ensuring ROM state set $\mathcal{X}^{i}$ - hard-coded
1:procedure GenerateROMstateSet(
ROM dimension $n$,
polynomial degree $i\in\mathbb{N}_{0}$
2:  if $i=1$ then return GenerateROMstateSet1(ROM dimension $n$)
3:  else if $i=2$ then return GenerateROMstateSet2(ROM dimension $n$)
4:  else if $i=3$ then return GenerateROMstateSet3(ROM dimension $n$)
5:  else if $i=8$ then return GenerateROMstateSet8(ROM dimension $n$)
6:  end if
7:end procedure
Appendix C ROM state errors with respect to the FOM
Figures 7, 7 and 8 show the relative average ROM state errors,
(81) $\frac{\left\|{\mathbf{X}}-{\mathbf{V}}_{n}{\mathbf{\tilde{X}}}\right\|_{F}}{\left\|{\mathbf{X}}\right\|_{F}},$
for Chafee-Infante equation in Section 4.1, ice sheet model in Section 4.2 and Burgers’ equation in Section 4.3, respectively. The matrix ${\mathbf{X}}=\left[{\bm{x}}_{0}\;\;\dots\;\;{\bm{x}}_{K_{\text{POD}}-1}\right]$ comprises the FOM snapshots used to construct the POD basis, and ${\mathbf{\tilde{X}}}=\left[\tilde{\bm{x}}_{0}\;\;\dots\;\;\tilde{\bm{x}}_{K_{\text{POD}}-1}\right]$ the corresponding ROM snapshots generated by integrating the ROM ODE (6) with the intrusive ROM operators and the inferred ROM operators, respectively. As can be seen, the ROM errors decrease monotonously for increasing ROM dimensions in all three test cases for both, intrusive and inferred ROMs. Additionally, we cannot see any difference between the errors of corresponding intrusive and inferred ROMs. In fact, the differences between these errors are all smaller than $10^{-10}$. This alignment is due to the exact reconstruction of the intrusive ROM operators shown in Fig. 2.
Figure 6: Relative average ROM state error (81) for Chafee-Infante equation.
Figure 7: Relative average ROM state error (81) for the ice sheet model.
Figure 8: Relative average ROM state error (81) for Burgers’ equation.